# GARG-AML against Smurfing: A Scalable and Interpretable Graph-Based Framework for Anti-Money Laundering

### v1

@@ Abstract
Money laundering poses a significant challenge as it is estimated to account for $2\%$$5\%$ of the global GDP. This has compelled regulators to impose stringent controls on financial institutions. One prominent laundering method for evading these controls, called smurfing, involves breaking up large transactions into smaller amounts. Given the complexity of smurfing schemes, which involve multiple transactions distributed among diverse parties, network analytics has become an important anti-money laundering tool. However, recent advances have focused predominantly on black-box network embedding methods, which has hindered their adoption in businesses. In this paper, we introduce GARG-AML, a novel graph-based method that quantifies smurfing risk through a single interpretable metric derived from the structure of the second-order transaction network of each individual node in the network. Unlike traditional methods, GARG-AML strikes an effective balance among computational efficiency, detection power and transparency, which enables its integration into existing AML workflows. To enhance its capabilities, we combine the GARG-AML score calculation with different tree-based methods and also incorporate the scores of the node’s neighbours. An experimental evaluation on large-scale synthetic and open-source networks demonstrate that the GARG-AML outperforms the current state-of-the-art smurfing detection methods. By leveraging only the adjacency matrix of the second-order neighbourhood and basic network features, this work highlights the potential of fundamental network properties towards advancing fraud detection.
@@ 1 Introduction
Money laundering disguises illegally-obtained funds, often from criminal activities, to make them appear legitimate. Once cleaned, the money can be used to finance further crimes. Estimates put the amount of laundered money at $2\%$ to $5\%$ of global GDP, which amounts to USD 2 trillion annually [42]. Since laundering money involves interactions with and through the normal financial system, regulators have imposed stringent anti-money laundering (AML) rules on financial institutions [16, 26].
Money laundering involves three main steps [42, 26]. Placement refers to the introduction of funds into the financial system. During layering, money is moved around to obscure its illegal origins. Finally, integration is the incorporation of laundered money into broader circulation. Because of the involvement of different steps and actors, network analytics has become an important tool for AML [11].
A popular money laundering method is smurfing [30, 46, 39, 51]. Large transactions between two accounts are often flagged according to business rules [6, 37, 34] in accordance with regulation [46]. Therefore, transactions are split into smaller amounts and transferred using multiple money mules—also called smurfs—making detection of illicit money flows harder since (1) the origin of the funds is obscured and (2) smaller transactions seem less suspicious.
Two types of patterns are typical for smurfing [39]: (1) scatter-gather and (2) gather-scatter [3]. Both patterns are visualised in Figure 1. With scatter-gather, money is sent to its destination account via money mules. Gather-scatter, on the other hand, consists of collecting money from multiple accounts into one (or a few) account(s) and sending it to many destination accounts. Gather-scatter is a popular money laundering method for cryptocurrencies [48].
Figure 1: The two smurfing patterns described in the literature [39, 3]; as scatter-gather (left) and gather-scatter (right).
To fight money laundering, AML practices use customer and transaction data to estimate each client’s money laundering propensity. When suspicion is raised, investigators build a suspicious activity report (SAR) to file with the regulator [16, 26]. Given the vast volume of monitored transactions and the involvement of different business departments in filing a SAR, two aspects are key to ensuring the effective business adoption of a method:
@@
Despite these business requirements, the current AML literature often focuses on deep graph neural networks (GNNs) [11]. GNNs are, however, not interpretable, prohibiting the adoption of these novel network analytics methods in practice. Consequently, many institutions still rely on business rules and manual feature engineering based on expert knowledge, which are expensive to maintain and easy to evade.
@@
  • We compare the current state-of-the-art smurfing detection methods, and we perform the first analysis of the scalability of these methods to very large transaction datasets.
The remainder of this work is organised as follows. Related work on anti-money laundering and smurfing is discussed in Section 2. The methodology behind the GARG-AML score calculation is described in Section 3. Section 4 outlines the experimental evaluation, and the results are presented in Section 5. We discuss the main limitations in Section 6, in terms of both the data in general and GARG-AML in particular. Section 7 concludes this work.
Complete implementations of the methods and experiments are provided on GitHub [^1] https://github.com/B-Deprez/GARG-AML to promote further adoption of and experimentation on the GARG-AML scores by peer researchers and practitioners.
@@ 2 Related Work
Anti-money laundering (AML), like other areas of financial fraud detection, traditionally relied on IF-THEN business rules—combinations of expert knowledge and regulatory requirements [6, 25, 36, 46, 7, 11]. These rules typically monitor individual transactions or aggregated amounts over time [16]. While offering transparency in decision-making [52], they are costly to maintain [43] and can be easily circumvented once exposed. As a result, machine learning has emerged as a preferred approach, offering scalability and the ability to detect complex laundering patterns [8].
@@
To capture interactions among people, network analytics for AML has been studied extensively in recent years [11]. Research has often relied on manual network feature engineering for profile definition. Dreżewski et al. [12] used different centrality measures, which resulted in nine roles in a criminal network. Fronzetti Colladon and Remondi [20] used similar metrics across industries and analysed the correlation with the target label to demonstrate the added value of network analytics for AML. Wu et al. [47] define temporal motives specifically tailored to money laundering and mixing detection in the Bitcoin network.
@@
In addition to manual feature engineering, graph neural networks (GNNs) have also been studied extensively [29, 2, 45, 11]. These methods capture the complex geometry of suspicious money laundering activities in financial networks by learning network topology while also incorporating node features. A drawback is that GNNs are not interpretable [22], which limits their use by investigators when filing their reports.
@@
Smurfing has long been a popular laundering method [46, 37]. Welling [46] suggested that it has been applied since 1970, when the Bank Secrecy Act was passed in the USA, which required reporting of transactions above \$$\$10,000$. To this day, similar rule-based systems are in place that require immediate investigation of transactions that surpass a given amount [37, 34, 6].
@@
Research has already been conducted on smurfing detection. Zhdanova et al. [51] introduced PSA@R for detecting smurfing in mobile money networks. It uses different underlying workflow models. Transaction streams that differ from the process model are flagged as anomalous. Additional criteria are applied to find money mules on the basis of the difference in the amounts of money coming into and going out of an account.
@@
FlowScope [27] tries to find anomalous sub-graphs on the basis of money flow. The authors constructed a k-partite network with two sets of nodes: nodes with net inflow to or net outflow from the bank and the remaining nodes, which correspond to the middle men. Li et al. [27] defined their anomaly score on the basis of the difference between the inflow and outflow of money at these middle men, and they claimed that smurfs transfer almost all the money they receive.
@@
A similar idea was put forward with AutoAudit [24]. The full adjacency matrix of the network is reordered, and depending on the density of the different blocks, a group of transactions is flagged as smurfing.
@@
Starnini et al. [39] presented a method that merges two transaction datasets and retains only those connections where both transactions occur within a specific time window. Prior to this merger, a substantial amount of data, i.e., nodes deemed trustworthy by the bank together with a majority of transactions, is discarded. After merging, the method retains certain motifs and applies an additional post-processing step, which then triggers manual investigations by AML experts.
@@
SMoTeF [38] is a three-step approach. The first is smurfing pattern extraction, using existing methods. The novelty comes from the second and third steps, where the patterns are checked against temporal constraints and maximal money flow, respectively. Many transactions are again pruned before making final predictions.
@@
Closely related to smurfing is mixing, which is a money laundering method that is used for cryptocurrencies [28]. Coins are transferred to the same entity and are redistributed afterward [48]. This obfuscates the history of ownership, which results in an inability to know the money’s real origin. Mixing can be viewed as the second smurfing pattern in Figure 1, i.e., gather-scatter.
@@
An anomaly score for mixing detection was introduced by Prado-Romero et al. [35]. The authors claimed that mixing services act as bridges between communities in a transaction network by linking otherwise unrelated users. After community detection, the anomaly score expresses the difference in a node’s intercommunity links compared with those of the other nodes in the same community.
@@
Wu et al. [47] constructed two networks for network feature extraction. The first links accounts using their transactions, whereas the second network represents both accounts and transactions as nodes to leverage transaction information. The extracted network features are based on predefined temporal motifs, which are added to the account and transaction features to detect smurfing in a positive and unlabelled (PU)-learning setting.
@@
The evaluation in most papers is done on relatively small datasets, making it harder to judge whether they scale well to the massive transaction networks faced by financial institutions. The dataset that was used in most of the above papers is the Czech dataset [1], which is a rather small network that contains approximately 11,000 nodes and approximately 273,000 edges. This dataset is unlabelled, so the authors of each paper inject synthetic pattern. Additionally, in most papers, evaluation and comparison is limited to threshold-dependent metrics like precision, recall and F1.
@@
Zhdanova et al. [51] Scatter-gather
Li et al. [27] Scatter-gather
Lee et al. [24] Scatter-gather
Starnini et al. [39] Scatter-gather & gather-scatter
Shadrooh and Nørvåg [38] Scatter-gather
Prado-Romero et al. [35] Gather-scatter
Wu et al. [47] Gather-scatter
@@
Zhdanova et al. [51] Simulated 10,000 Precision, recall
Li et al. [27] Proprietary & Czech Financial 6,130,000 & 11,374 F1, Time
Lee et al. [24] Czech Financial & Enron 11,374 & 16,771 Accuracy
Starnini et al. [39] Proprietary 18,000,000 (reduced to 21,000) Time
Shadrooh and Nørvåg [38] Czech Financial & Enron & Darpa 11,374 & 16,771 & 371,000 Time, Precision, Recall, F1
Prado-Romero et al. [35] Bitcoin 942,204 Recall
Wu et al. [47] WalletExplorer ±$\pm 2,500,000$ TPR, FPR, geometric mean
@@ 3.1 Preliminaries and Notation
Suppose $\mathcal{N}_{k}(v)$ represents the set of nodes that are at distance exactly $k$ from node $v$. For GARG-AML, we consider the neighbours of each node $v$ at distances one and two, i.e., $\mathcal{N}_{1}(v)$ and $\mathcal{N}_{2}(v)$, respectively. Suppose that $\mid\mathcal{N}_{1}(v)\mid=n$ and $\mid\mathcal{N}_{2}(v)\mid=m-1$; then, the adjacency matrix of the second-order neighbourhood of $v$ is of size $(m+n)\times(m+n)$.
@@ 3.2 Undirected
We introduce GARG-AML using the smurfing pattern defined by Starnini et al. [39] and illustrated in Figure 2. The full GARG-AML method is given in Algorithm 1. Given node $A$, the corresponding adjacency matrix - defined in Equation (1) - is constructed by arranging the nodes in a specific order: node $A$ is placed first, followed by the second-order neighbour $E$, and then its first-order neighbours $B,C$ and $D$. This ordering results in an adjacency matrix that is divided into four distinct blocks. The diagonal blocks contain only zeros, whereas the off-diagonal blocks contain ones.
In this example, we will consider only the gather-scatter pattern, as this is most widely recognised as smurfing. The GARG-AML calculations and node order stay the same in the gather-scatter case.
Figure 2: Example of a smurfing pattern. $A$ is the node that will receive the score. The first-order neighbours ($B,C,D$) are shown in purple, and the second-order neighbour ($E$) is shown in green.
@@
In general, the first row of the adjacency matrix, which corresponds to entries $(A_{1,j})$, represents the connections of node $v$ under consideration. The next $m-1$ rows, $(A_{i,j})$ for ${2\leq i\leq m}$, represent the connections of nodes that are located at a distance of two from $v$. The final $n$ rows, entries $(A_{i,j})$ for ${m+1\leq i\leq m+n}$, represent the connections of nodes at a distance of one from $v$.
By enforcing this node ordering—the target node, second-order neighbours, and first-order neighbours—for the adjacency matrix, we ensure that when a pure smurfing pattern is present, the diagonal blocks remain sparse (low density), and the off-diagonal blocks have a high density. The exact order of nodes within each group is not important. For example, any permutation within the first-order neighbours, such as $C,B,D$, is equally valid.
@@
GARG-AML measures the similarity between an observed pattern and the pure smurfing template by analysing the expected density and actual density (or sparsity) of each block. This results in a single quantitative score expressing similarity to the pure smurfing pattern.
In computing these densities, several structural properties of the adjacency matrix are taken into account. First, self-loops are excluded; thus, diagonal elements are always zero and are omitted from the sparsity/density calculations. Furthermore, in $\text{block}_{1}$ (upper-left block), the first row and column can never contain ones, as a ‘1’ would indicate a direct connection between a node and its second-order neighbours—a contradiction. This block serves as a check that the nodes at distance two do no directly make transactions to each other, in case multiple nodes at a distance of two are present. Similarly, for $\text{block}_{2}$ (upper-right block), all the elements in the first row must be ones since these represent connections between the central node and its first-order neighbours.
These structural constraints affect both the numerator and denominator in the density calculations. Accordingly, when computing block-wise densities, we consider only the valid entries, and the block sizes (denominators) are normalised to reflect this. On the basis of the observations above, smurfing is likely to be present when the density of $\text{block}_{2}$ (upper-right block) is high, while $\text{block}_{1}$ (upper-left block) and $\text{block}_{3}$ (lower-right) exhibit densities close to zero. This is reflected in the final score as follows:
@@
$\text{score}_{\text{total}}$ $=$ $\text{score}_{2}-\frac{l_{1}\cdot\text{score}_{1}+l_{3}\cdot\text{score}_{3}}{l_{1}+l_{3}}$ (8)
The inclusion of $l_{i}$ increases the flexibility of the score calculation. If $l_{i}=1$, the score is the average of $\text{score}_{1}$ and $\text{score}_{3}$. In our experiments, however, we take the weighted average by setting:
@@
When there is only one destination account, there is only one node at distance two (as in Figure 2), which results in $l_{1}$ being 0. We take the weighted average because we need to compare only $\text{score}_{2}$ to $\text{score}_{3}$.
The final score in Equation (8) ranges from $-1$ to $1$. When the second-order neighbourhood resembles a smurfing pattern, the final score is close to 1.

### v2

@@ Abstract
Purpose: This paper introduces a novel graph-based method, GARG-AML, for efficient and effective anti-money laundering (AML). It quantifies smurfing risk, a popular money laundering method, by providing each node in the network with a single interpretable score. The proposed method strikes a balance among computational efficiency, detection power and transparency. Different versions of GARG-AML are introduced for undirected and directed networks.
Methodology: GARG-AML constructs the adjacency matrix of a node’s second-order neighbourhood in a specific way. This allows us to use the density of different blocks in the adjacency matrix to express the neighbourhood’s resemblance to a pure smurfing pattern. GARG-AML is extended using a decision tree and gradient-boosting classifier to increase its performance even more. The methods are tested on synthetic and on open-source data against the current state-of-the-art in AML.
Findings: We find that GARG-AML obtains state-of-the-art performance on all datasets. We illustrate that GARG-AML scales well to massive transactions graphs encountered at financial institutions. By leveraging only the adjacency matrix of the second-order neighbourhood and basic network features, this work highlights the potential of fundamental network properties towards advancing fraud detection.
Originality: This paper uses only basic network features and expert knowledge on smurfing to construct a performant AML system. The originality lies in the translation of smurfing detection to these features and network representation. Our proposed method is built around the real business needs of scalability and interpretability. It therefore provides a solution that can be easily implemented at financial institutions or incorporated in existing AML solutions.
@@ 1 Introduction
Money laundering disguises illegally-obtained funds, often from criminal activities, to make them appear legitimate. Once cleaned, the money can be used to finance further crimes. Estimates put the amount of laundered money at $2\%$ to $5\%$ of global GDP, which amounts to USD 2 trillion annually (United Nations Office on Drugs and Crime, (n.d.). Since laundering money involves interactions with and through the normal financial system, regulators have imposed stringent anti-money laundering (AML) rules on financial institutions (FATF, 2012-2023; Levi & Reuter, 2006).
Money laundering involves three main steps (United Nations Office on Drugs and Crime, (n.d.; Levi & Reuter, 2006). Placement refers to the introduction of funds into the financial system. During layering, money is moved around to obscure its illegal origins. Finally, integration is the incorporation of laundered money into broader circulation. Because of the involvement of different steps and actors, network analytics has become an important tool for AML (Deprez et al., 2024).
A popular money laundering method is smurfing (Madinger, 2012; Welling, 1989; Starnini et al., 2021; Zhdanova et al., 2014). Large transactions between two accounts are often flagged according to business rules (Bolton & Hand, 2002; Senator et al., 1995; Oztas et al., 2024) in accordance with regulation (Welling, 1989). Therefore, transactions are split into smaller amounts and transferred using multiple money mules—also called smurfs—making detection of illicit money flows harder since (1) the origin of the funds is obscured and (2) smaller transactions seem less suspicious.
Two types of patterns are typical for smurfing (Starnini et al., 2021): (1) scatter-gather and (2) gather-scatter (Altman et al., 2023). Both patterns are visualised in Figure 1. With scatter-gather, money is sent to its destination account via money mules. Gather-scatter, on the other hand, consists of collecting money from multiple accounts into one (or a few) account(s) and sending it to many destination accounts. Gather-scatter is a popular money laundering method for cryptocurrencies (Wu, Hu, Zhou, Wang, Luo, Wang, Zhang & Ren, 2021).
Figure 1: The two smurfing patterns described in the literature (Starnini et al., 2021; Altman et al., 2023); as scatter-gather (left) and gather-scatter (right).
To fight money laundering, AML practices use customer and transaction data to estimate each client’s money laundering propensity. When suspicion is raised, investigators build a suspicious activity report (SAR) to file with the regulator (FATF, 2012-2023; Levi & Reuter, 2006). Given the vast volume of monitored transactions and the involvement of different business departments in filing a SAR, two aspects are key to ensuring the effective business adoption of a method:
@@
Despite these business requirements, the current AML literature often focuses on deep graph neural networks (GNNs) (Deprez et al., 2024). GNNs are, however, not interpretable, prohibiting the adoption of these novel network analytics methods in practice. Consequently, many institutions still rely on business rules and manual feature engineering based on expert knowledge, which are expensive to maintain and easy to evade.
@@
  • We compare the current state-of-the-art smurfing detection methods, and we perform the first analysis of the scalability of these methods to very large transaction datasets;
  • Complete implementations of the methods and experiments are provided on GitHub to promote further adoption of and experimentation on the GARG-AML scores by peer researchers and practitioners.
The remainder of this work is organised as follows. Related work on anti-money laundering and smurfing is discussed in Section 2. The methodology behind the GARG-AML score calculation is described in Section 3. Section 4 outlines the experimental evaluation, and the results are presented in Section 5. We discuss the main limitations, in terms of both the data in general and GARG-AML in particular, in Section 6. Section 7 concludes this work.
@@ 2 Related Work
Anti-money laundering (AML), like other areas of financial fraud detection, traditionally relied on IF-THEN business rules—combinations of expert knowledge and regulatory requirements (Bolton & Hand, 2002; Leonard, 1995; Sánchez et al., 2009; Welling, 1989; Cardoso et al., 2022; Deprez et al., 2024). These rules typically monitor individual transactions or aggregated amounts over time (FATF, 2012-2023). While offering transparency in decision-making (Zhu & Hu, 2013), they are costly to maintain (Van Belle et al., 2023) and can be easily circumvented once exposed. As a result, machine learning has emerged as a preferred approach, offering scalability and the ability to detect complex laundering patterns (Chen et al., 2018).
@@
To capture interactions among people, network analytics for AML has been studied extensively in recent years (Deprez et al., 2024). Research has often relied on manual network feature engineering for profile definition. Dreżewski et al. (2015) used different centrality measures, which resulted in nine roles in a criminal network. Fronzetti Colladon & Remondi (2017) used similar metrics across industries and analysed the correlation with the target label to demonstrate the added value of network analytics for AML. Wu, Liu, Chen, Huang, Zheng & Zhang (2021) define temporal motives specifically tailored to money laundering and mixing detection in the Bitcoin network.
@@
In addition to manual feature engineering, graph neural networks (GNNs) have also been studied extensively (Lo et al., 2023; Alarab et al., 2020; Weber et al., 2019; Deprez et al., 2024). These methods capture the complex geometry of suspicious money laundering activities in financial networks by learning network topology while also incorporating node features. A drawback is that GNNs are not interpretable (Kute et al., 2021), which limits their use by investigators when filing their reports.
@@
Smurfing has long been a popular laundering method (Welling, 1989; Senator et al., 1995). Welling (1989) suggested that it has been applied since 1970, when the Bank Secrecy Act was passed in the USA, which required reporting of transactions above $\mathdollar 10,000$. To this day, similar rule-based systems are in place that require immediate investigation of transactions that surpass a given amount (Senator et al., 1995; Oztas et al., 2024; Bolton & Hand, 2002).
@@
Research has already been conducted on smurfing detection. Zhdanova et al. (2014) introduced PSA@R for detecting smurfing in mobile money networks. It uses different underlying workflow models. Transaction streams that differ from the process model are flagged as anomalous. Additional criteria are applied to find money mules on the basis of the difference in the amounts of money coming into and going out of an account.
@@
FlowScope (Li et al., 2020) tries to find anomalous sub-graphs on the basis of money flow. The authors constructed a k-partite network with two sets of nodes: nodes with net inflow to or net outflow from the bank and the remaining nodes, which correspond to the middle men. Li et al. (2020) defined their anomaly score on the basis of the difference between the inflow and outflow of money at these middle men, and they claimed that smurfs transfer almost all the money they receive.
@@
A similar idea was put forward with AutoAudit (Lee et al., 2020). The full adjacency matrix of the network is reordered, and depending on the density of the different blocks, a group of transactions is flagged as smurfing.
@@
Starnini et al. (2021) presented a method that merges two transaction datasets and retains only those connections where both transactions occur within a specific time window. Prior to this merger, a substantial amount of data, i.e., nodes deemed trustworthy by the bank together with a majority of transactions, is discarded. After merging, the method retains certain motifs and applies an additional post-processing step, which then triggers manual investigations by AML experts.
@@
SMoTeF (Shadrooh & Nørvåg, 2024) is a three-step approach. The first is smurfing pattern extraction, using existing methods. The novelty comes from the second and third steps, where the patterns are checked against temporal constraints and maximal money flow, respectively. Many transactions are again pruned before making final predictions.
[…]
@@
Closely related to smurfing is mixing, which is a money laundering method that is used for cryptocurrencies (Liu et al., 2021). Coins are transferred to the same entity and are redistributed afterward (Wu, Hu, Zhou, Wang, Luo, Wang, Zhang & Ren, 2021). This obfuscates the history of ownership, which results in an inability to know the money’s real origin. Mixing can be viewed as the second smurfing pattern in Figure 1, i.e., gather-scatter.
@@
An anomaly score for mixing detection was introduced by Prado-Romero et al. (2018). The authors claimed that mixing services act as bridges between communities in a transaction network by linking otherwise unrelated users. After community detection, the anomaly score expresses the difference in a node’s intercommunity links compared with those of the other nodes in the same community.
@@
Wu, Liu, Chen, Huang, Zheng & Zhang (2021) constructed two networks for network feature extraction. The first links accounts using their transactions, whereas the second network represents both accounts and transactions as nodes to leverage transaction information. The extracted network features are based on predefined temporal motifs, which are added to the account and transaction features to detect smurfing in a positive and unlabelled (PU)-learning setting.
@@
The evaluation in most papers is done on relatively small datasets, making it harder to judge whether they scale well to the massive transaction networks faced by financial institutions. The dataset that was used in most of the above papers is the Czech dataset (Data World, 1999), which is a rather small network that contains approximately 11,000 nodes and approximately 273,000 edges. This dataset is unlabelled, so the authors of each paper inject synthetic pattern. Additionally, in most papers, evaluation and comparison is limited to threshold-dependent metrics like precision, recall and F1.
@@
Zhdanova et al. (2014) Scatter-gather
Li et al. (2020) Scatter-gather
Lee et al. (2020) Scatter-gather
Starnini et al. (2021) Scatter-gather & gather-scatter
Shadrooh & Nørvåg (2024) Scatter-gather
Prado-Romero et al. (2018) Gather-scatter
Wu, Liu, Chen, Huang, Zheng & Zhang (2021) Gather-scatter
@@
Zhdanova et al. (2014) Simulated 10,000 Precision, recall
Li et al. (2020) Proprietary & Czech Financial 6,130,000 & 11,374 F1, Time
Lee et al. (2020) Czech Financial & Enron 11,374 & 16,771 Accuracy
Starnini et al. (2021) Proprietary 18,000,000 (reduced to 21,000) Time
Shadrooh & Nørvåg (2024) Czech Financial & Enron & Darpa 11,374 & 16,771 & 371,000 Time, Precision, Recall, F1
Prado-Romero et al. (2018) Bitcoin 942,204 Recall
Wu, Liu, Chen, Huang, Zheng & Zhang (2021) WalletExplorer $\pm 2,500,000$ TPR, FPR, geometric mean
@@ 3 GARG-AML
The main idea of GARG-AML is to provide each individual node with a score expressing how likely it is that this entity is involved in a smurfing scheme. This is done by analysing key characteristics of the second-order neighbourhood of the nodes. Below, we start by formally defining what we mean with second-order neighbourhood. Afterwards, we provide two version for GARG-AML, one for undirected and one for directed networks. The full GARG-AML method is given in Algorithm 1 in the experimental set-up in Section 4.2.
@@ 3.1 Preliminaries and Notation
Suppose $\mathcal{N}_{k}(v)$ represents the set of nodes that are at distance exactly $k$ from node $v$, with $\mathcal{N}_{0}(v)=\{v\}$. For GARG-AML, we consider the neighbours of each node $v$ at distances one and two, i.e., $\mathcal{N}_{1}(v)$ and $\mathcal{N}_{2}(v)$, respectively. Suppose that $\mid\mathcal{N}_{1}(v)\mid=n$ and $\mid\mathcal{N}_{2}(v)\mid=m-1$; then, the adjacency matrix of the second-order neighbourhood of $v$ is of size $(m+n)\times(m+n)$.
The $k$th-order neighbourhood of node $v$ is the sub-network $G(V^{\prime},E^{\prime})$ where $V^{\prime}=\bigcup\limits_{l\leq k}\mathcal{N}_{k}(v)$ and $E^{\prime}=\{(u,w)\mid u,w\in V^{\prime}\vee(u,w)\in E\}$. Hence, it is the sub-network that representing all nodes at distance at most $k$ and all the interaction among these nodes.
@@ 3.2 Undirected
We introduce GARG-AML using the smurfing pattern defined by Starnini et al. (2021) and illustrated in Figure 2. In this example, we want to give node $A$ a score. The adjacency matrix of its second-order neighbourhood, given in Equation (1), is constructed by arranging the nodes in a specific order: node $A$ is placed first, followed by the second-order neighbour $E$, and then its first-order neighbours $B,C$ and $D$. This ordering results in an adjacency matrix that is divided into four distinct blocks. The diagonal blocks contain only zeros, whereas the off-diagonal blocks contain ones.
In this example, we will consider only the scatter-gather pattern, as this is most widely recognised as smurfing. The GARG-AML calculations and node order stay the same in the gather-scatter case.
Figure 2: Example of a smurfing pattern. AA is the node that will receive the score. The first-order neighbours ($B,C,D$) are shown in purple, and the second-order neighbour ($E$) is shown in green.
@@
Given the sights of the example above, we define the general GARG-AML method, starting with the adjacency matrix. The first row of the adjacency matrix, which corresponds to entries $(A_{1,j})$, represents the connections of node $v$ under consideration. The next m1m-1 rows, i.e., $(A_{i,j})$ for $2\leq i\leq m$}, represent the connections of nodes that are located at a distance of two from $v$. The final $n$ rows, i.e., $(A_{i,j})$ for $m+1\leq i\leq m+n$}, represent the connections of nodes at a distance of one from $v$.
Enforcing this node ordering—the target node, second-order neighbours, and first-order neighbours—for the adjacency matrix ensures that the diagonal blocks remain sparse (low density), and the off-diagonal blocks have a high density in case we are faced with a pure smurfing pattern. The exact order of nodes within each group is not important. For example, any permutation within the first-order neighbours, such as $C,B,D$, is equally valid.
@@
The density, i.e., the relative number of non-zero elements in each block are essential for our method. GARG-AML measures the similarity between an observed pattern and the pure smurfing template by analysing the expected density and actual density (or sparsity) of each block. This results in a single quantitative score expressing similarity to the pure smurfing pattern.
Several structural properties of the adjacency matrix are taken into account to come to more rigorous density calculations. First, self-loops are excluded; thus, diagonal elements are always zero and are not considered from the sparsity/density calculations. Furthermore, in $\text{block}_{1}$ (upper-left block), the first row and column can never contain ones, as a ‘1’ would indicate a direct connection between a node and its neighbours at distance two—a contradiction. This upper-left block serves as a check that the nodes at distance two do no directly make transactions to each other, in case multiple nodes at a distance of two are present. Similarly, for $\text{block}_{2}$ (upper-right block), all the elements in the first row must be ones since these represent connections between the node under investigation and its first-order neighbours.
We consider only the free entries, i.e., those that are not certain upfront, for density calculations. For readability, we define the block size as the number of free elements. This is reflected in both the numerator (number of non-zero elements) and denominator (block size).
@@
Smurfing is likely to be present when the density of $\text{block}_{2}$ (upper-right block) is high, while $\text{block}_{1}$ (upper-left block) and $\text{block}_{3}$ (lower-right) exhibit densities close to zero. Therefore, the final GARG-AML score is calculated as=:
$\text{score}_{\text{total}}=\text{score}_{2}-\frac{l_{1}\cdot\text{score}_{1}+l_{3}\cdot\text{score}_{3}}{l_{1}+l_{3}},$ (8)
where $l_{i}$ is a weight given to each term in the score calculation. The inclusion of $l_{i}$ increases the flexibility of the score calculation. If $l_{i}=1$, the score is the average of $\text{score}_{1}$ and $\text{score}_{3}$. In our experiments, however, we take a weighted average using the block sizes:
@@
When there is only one destination account, there is only one node at distance two (as in Figure 2). This results in $l_{1}$ being 0. Then, the calculations simplify to $\text{score}_{\text{total}}=\text{score}_{2}-\text{score}_{3}$, and we only analyse whether all money ends up at the destination node (high $\text{score}_{2}$) and none of the intermediary accounts interact with each other (low $\text{score}_{3}$).
The final score in Equation (8) ranges from 1-1 to $1$. When the second-order neighbourhood resembles a smurfing pattern, the final score is close to 1. To illustrate how these calculations are done, additional examples are provided in Appendix A.1.